OpenStudy (anonymous):

Easy exlaplanation how to work these problems PLEASE. . . . #1. . 34-{69-4[2+3(2^2)]} #2. . .18 divided by 3+{2 times 9-[6+(2+1]}

OpenStudy (anonymous):

34-(69-4(2+3(2^2))) = 34-(69-4(2+3(4))) = 34-(69-4(2+12)) = 34-(69-4(14)) = 34-(69-56)) = 34 - 13 = 21

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

:)